OpenStudy (anonymous):

WHAT WOULD I DO FOR THIS ONE A small rock sits on the edge of a tall building. A strong wind blows the rock off the edge. The distance, in feet, between the rock and the ground t seconds after the rock leaves the edge is given by d = −16t 2 − 4t + 420. How long after the rock leaves the edge is it 390 feet from the ground?

OpenStudy (anonymous):

set = 390 and solve for t

OpenStudy (anonymous):

put d = 390 and solve fr t

OpenStudy (anonymous):

\[-16t^2-4t+420=390\] \[-16t^2-4t+60=0\] etc

OpenStudy (anonymous):

so 390=-16t^2-4t+420?

OpenStudy (anonymous):

ok thanks!

OpenStudy (anonymous):

except that my arithmetic is wrong. it is \[-16t^2-4t+420=390\] \[-16t^2-4t+30=0\]

OpenStudy (anonymous):

i get 5/4